OpenStudy (anonymous):

WILL MEDAL + FAN!!! The product of a number and -4 is at least the difference of 21 and the number. Find the greatest valid real number.

OpenStudy (anonymous):

AGEEN

OpenStudy (anonymous):

ill help

OpenStudy (emmamink):

-4n\[-4n \ge 21 - n\]

OpenStudy (anonymous):

-4n −4n≥21−n

OpenStudy (emmamink):

ok, so now we can add n to each side to get 21 by itself.\[-3n \ge 21\]

OpenStudy (emmamink):

Now, I think you can figure it out from here!

OpenStudy (anonymous):

can someboddy fan and medal me i am very sad ;(

OpenStudy (emmamink):

Divide each side by -3 and we get our answer! \[n \ge -7\]

OpenStudy (emmamink):

yes!

OpenStudy (emmamink):

Thanks 4 the medals everyone!